# Engineered Josephson diode effect in kinked Rashba nanochannels

### v1

@@ I Introduction
A superconducting diode effect (SDE) refers to a phenomenon observed in certain superconducting devices where the flow of electrical current is significantly influenced by the direction of the current biasing the junction [1]. It essentially enables supercurrent to flow preferentially in one direction, providing new functionalities for superconducting circuits, such as superconducting quantum interference devices (SQUIDs), superconducting magnetic energy storage systems, and superconducting digital electronics. In recent years, there has been a significant progress towards the realization of the diode effect in superconducting heterostructures [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. This phenomenon has also been investigated in several Josephson junction (JJ) systems, reflecting the advancement from pure theoretical studies [15, 16, 17, 18, 19, 20, 21, 22, 23] to tangible implementations. The latters exploit different configurations, and extend to the inclusion of unconventional superconductors [24, 25, 26, 27, 28, 29], graphene [30, 31], and topological materials [32, 33, 34, 35, 36]. The effectiveness of these proposals is typically evaluated through the SDE efficiency, which is determined by the difference of the Josephson critical currents when the junction is biased in opposite directions normalized to the sum of these currents. It has been explicitly demonstrated how nonreciprocity of supercurrent is intertwined with underlying symmetries of the system. In fact, nonreciprocity driven by breaking of space-inversion symmetry and time reversal symmetry, in the presence, e.g., of Rashba spin-orbit interaction (RSOC) and an applied magnetic field, has been widely investigated. In particular, evidences of nonreciprocity have been recently observed in the magnetic field dependence of the critical current in JJs at oxides interfaces, which are known to behave as Rashba-like systems [37, 38]. An asymmetry in the critical current as a function of the applied magnetic field, $I_{c}(H)\neq I_{c}(-H)$, has been revealed, alongside an anomalous amplification of the critical current attributed to the emergence of Majorana bound states (MBSs) [39].
@@
On the other side, the interplay between RSOC and geometric deformation of the JJ can also generate intriguing effects on the Josephson current, inducing an anomalous phase shift in the ground state of the junction and favouring the nonreciprocity of supercurrent [40]. The latter originates from non-abelian phases controlled by the RSOC [41], impacting on the overall transport properties of the system. Geometric control of Josephson coupling has been also recently demonstrated in twisted cuprate van der Waals heterostructures [42, 43, 44] and the emergence of SDE in artificial junctions composed of twisted layers of Bi2Sr2CaCu2222O8+δ [45, 46] has been shown. These evidences suggest that a rich phenomenology could emerge for the non-reciprocal transport of supercurrent in the presence of RSOC, modulated by the geometric deformation of the junction and in the vicinity of topological phases. Understanding and harnessing this interplay hold great promise for the development of superconducting diode devices with enhanced functionalities, ranging from efficient energy conversion to robust quantum information processing.
@@
To this aim, we analyze the diode effect in a short superconductor-normal-superconductor (SNS) junction with a kinked nanostrip geometry in the presence of strong Rashba spin-orbit coupling (RSOC).
@@
We show a geometrically controlled SDE by changing the kink angle $\theta$ and map the diode efficiency by scanning the angle and the external magnetic field perpendicular to the RSOC plane. Interestingly the maximal efficiency is obtained on the verge of the topological phase where non-trivial zero-energy bound states emerge. Beyond the non-reciprocal behavior of the supercurrent, we demonstrate the evolution of Andreev bound states at the Fermi level from trivial to non-trivial Majorana bound states (MBSs) with unique spectral characteristics. The organization of the paper is as follows: In Sec.II, we introduce the model and formalism. In Sec. III, we discuss the rich behavior of the supercurrent as a function of the kink angle and applied magnetic field, focusing on the supercurrent diode effect (SDE). Finally, in Sec. IV, we present our conclusions and provide an outlook.
@@ II Model and formalism
We consider a kinked nanostrip superconducting junction on a two-dimensional gas with Rashba interaction and variable kink angle. The superconducting leads are considered at a relative angle $\theta$ with respect to the positive orientation of the $x$-axis of the square lattice (see Fig. 1). The region defined by the junction determines the system domain $\mathcal{D}$, with its area given by $A=L_{W}(2L_{S}+L_{N})$. The lengths of the superconducting and normal regions projected along the $x$-axis are denoted by $L_{S}$ and $L_{N}$ respectively, while $L_{W}$ represents the width of the junction (see supplemental materials at VI).
Figure 2: (a) Phase diagram of the critical current as a function of $\theta$ and $M_{z}$, the white lines correspond to $\theta_{n}=\arctan(3/n)$ with $n=4,6,8,10,12,18,24$. (b) Selected horizontal cuts from panel (a) showing a diode effect strongly marked for high values of $\theta$. (c) Phase diagram of the diode efficiency $\eta$ and (d) field-optimization of $\eta$ as a function of $\theta$. The superimposed orange line depicts the moving average computed over a 25-point data window. (e) Phase diagram of the zero-phase current behavior, with select horizontal cuts provided in panel (f). Notably, red regions correspond to the nucleation of $\pi/2$-junctions, easily recognizable by the enhanced regions in the green curves of the panel (f) ($\theta=0.4$, $0.7$). The $\theta$ values chosen in panels (b) and (f) are also signaled in the phase diagrams (a) and (e) by means of horizontal dashed lines.
@@ III Results
Figure 3: Classification of the sub-gap states, ABSs, MBSs and ZABSs, by means of the CPRs (a), (d), (g), the probability densities $|\psi_{i}^{2}|$ (b), (e), (h) and spectra and DOSs (c), (f), (i). The graphs and the DOSs are calculated respectively by fixing $\phi$ at the value marked by the red point in the CPRs and the vertical red cut in the spectra. The graphs reproduce the geometry of a nano-strip with $L_{S}=100$, $L_{N}=3$, $L_{W}=3$, being the radius of the circles proportional to $|\psi_{i}^{2}|$ ($i=1,\dots,L$). The spectral energy $E$ and the Majorana polarization $P_{M}(E)$ of these states are also reported.
@@
Apart from MBSs, the peculiar intertwining between the effects controlled by the geometry (kink angle) and the magnetic energy engenders additional sub-gap quasiparticle modes. We classify all the sub-gap states into three categories, distinguished by their spectral properties and influence on the CPR (Fig. 3). Trivial Andreev bound states (ABSs) are associated with sinusoidal CPRs (Fig. 3(a)) and their probability densities are spread over the whole system size (Fig. 3(b)), carrying a sub-gap excitation energy well above the Fermi level (Fig. 3(g)). These states can be found in the region where $I_{c}$ decreases with respect to $M_{z}$. MBSs, on the other hand, are associated with an enhancement of $I_{c}$ vs $M_{z}$ curves and exhibit a sawtooth CPR with a single nodal point in the phase-periodicity range (Fig. 3(d)). They are localized at the four edges of the junction (Fig. 3(e)) and reside at the Fermi energy (Fig. 3(f)) [65, 66]. Finally, Zero-energy Andreev bound states (ZABSs) can be found in the region where $I_{c}\sim I_{c}(M_{z}=0)$, before the critical current increases. The phase-space region, defined by $(\theta,M_{z})$ and hosting ZABSs, exhibits complex boundaries, the prediction of which is a rather complex task. Concerning their influence on the transport properties, ZABSs exhibit a double-sawtooth CPR (Fig. 3(g)) and, compared to MBSs, exhibit localization on a smaller number of lattice sites, placed at the left corner of the superconducting leads (Fig. 3(h)). They also share the same energy of MBSs for two specific values of the superconducting phase difference (Fig. 3(i)). However, despite their similarities to MBSs, neither ABSs nor ZABSs carry topological charge, as evidenced by the Majorana polarization ($P_{M}$[67, 68], which is indicated in the insets of panels (b), (e), and (f). Indeed, the Majorana polarization, a measure of topological quasiparticle weight in the Nambu space, is equal to $1$ for genuine MBSs, while $P_{M}\ll 1$ for ABSs and ZABSs. Moreover, ZABs can be understood as defect states dressed by the superconducting pairing, which originates from connectivity defects or quasi-periodicity of the lattice structure [69, 70].
@@ IV Conclusions and outlooks
In summary, we have reported the emergence of the superconducting diode effect in geometry-controlled Josephson junctions based on Rashba nanostrips. By tuning parameters such as the magnetic field and the kink angle, $M_{Z}$ and $\theta$, respectively, we achieved control over critical current and diode efficiency, which is pivotal for device optimization. Notably, we achieved remarkable optimization of diode efficiency, with values exceeding $40\%$ and a minimum efficiency of $20\%$ for $\theta\geq 0.1$. The intricate interplay between effects controlled by $\theta$ and $M_{Z}$ reveals additional intriguing phenomena, including the crossover from conventional to anomalous Josephson effects and the effective tuning of the quasiparticle spectrum. Our investigation identifies Andreev bound states, zero-energy Andreev bound states, and Majorana bound states as significant sub-gap excitations, each associated with distinct transport regimes of the junction. These phases are highly sensitive to the nanochannel bending, thus providing a relevant resource to optimize the SDE. Also for real systems, for which the fine geometrical tuning may not be achieved with a large enough resolution, we demonstrated the robustness of the SDE by increasing the kink angle, i.e., the higher $\theta$, the larger the SDE. Moreover, we cannot exclude that actual physical systems do not present asymmetries, angular misalignments, and/or irregularities at the edges. Indeed, the presence of impurities has been demonstrated to induce transition to $\pi$-junctions in spin-orbit magnetic devices [71, 72]. This aspect suggests the potential coexistence of different phases in a realistic experimental setup close to specific junction configurations, which can be further investigated. In conclusion, our findings represent a significant advancement in both the fundamental understanding and practical engineering of Rashba-based Josephson junctions for future nanoelectronic devices.
@@ VI.3 Further results on the anomalous Josephson effect
In Fig. 5(a) we consider a horizontal cut extracted from the phase diagram depicted in Fig. 2(b), with $\theta=0.2$. A dip-peak structure is evident around $M_{z}\sim 0.055$ (vertical dashed red line), marking the critical threshold for the onset of topological phases. We observe that the spectral behavior, initially lifted from zero energy at lower fields with a sinusoidal CPR (see Fig. 5(a)), converges towards zero energy for $M_{z}\geq 0.055$ (see Figs. 5(b)-(c)), accompanied by the emergence of a sawtooth CPR in Fig. 5(c).

### v2

@@ I Introduction
A superconducting diode effect (SDE) refers to a phenomenon observed in certain superconducting devices where the flow of electrical current is significantly influenced by the direction of the current biasing the junction [1]. It essentially enables supercurrent to flow preferentially in one direction, providing new functionalities for superconducting circuits, such as superconducting quantum interference devices (SQUIDs), superconducting magnetic energy storage systems, and superconducting digital electronics. In recent years, there has been a significant progress towards the realization of the diode effect in superconducting heterostructures [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. This phenomenon has also been investigated in several Josephson junction (JJ) systems, reflecting the advancement from pure theoretical studies [15, 16, 17, 18, 19, 20, 21, 22, 23] to tangible implementations. The latters exploit different configurations, and extend to the inclusion of unconventional superconductors [24, 25, 26, 27, 28, 29], graphene [30, 31], and topological materials [32, 33, 34, 35, 36]. The effectiveness of these proposals is typically evaluated through the SDE efficiency, which is determined by the difference of the Josephson critical currents when the junction is biased in opposite directions normalized to the sum of these currents. It has been explicitly demonstrated how nonreciprocity of supercurrent is intertwined with underlying symmetries of the system. In fact, nonreciprocity driven by breaking of space-inversion symmetry and time reversal symmetry, in the presence, e.g., of Rashba spin-orbit interaction (RSOC) and an applied magnetic field, has been widely investigated. In particular, evidences of nonreciprocity have been recently observed in the magnetic field dependence of the critical current in JJs at oxides interfaces, which are known to behave as Rashba-like systems [37, 38]. An asymmetry in the critical current as a function of the applied magnetic field, $I_{c}(H)\neq I_{c}(-H)$, has been revealed, alongside an anomalous amplification of the critical current attributed to the emergence of Majorana bound states (MBSs) [39, 40].
@@
On the other side, the interplay between RSOC and geometric deformation of the JJ can also generate intriguing effects on the Josephson current, inducing an anomalous phase shift in the ground state of the junction [41, 42] and favouring the nonreciprocity of supercurrent [43, 44, 45, 46]. The latter originates from non-abelian phases controlled by the RSOC [47], impacting on the overall transport properties of the system. Geometric control of Josephson coupling has been also recently demonstrated in twisted cuprate van der Waals heterostructures [48, 49, 50] and the emergence of SDE in artificial junctions composed of twisted layers of Bi2Sr2CaCu2222O8+δ [51, 52] has been shown. These evidences suggest that a rich phenomenology could emerge for the non-reciprocal transport of supercurrent in the presence of RSOC, modulated by the geometric deformation of the junction and in the vicinity of topological phases. Understanding and harnessing this interplay hold great promise for the development of superconducting diode devices with enhanced functionalities, ranging from efficient energy conversion to robust quantum information processing.
@@
To this aim, we analyze the diode effect in a short superconductor-normal-superconductor (SNS) junction with a kinked nanostrip geometry in the presence of strong Rashba spin-orbit coupling (RSOC). The kink angle of the Josephson junction affects transport properties by modifying the energy profile and altering the conductance. It influences the critical current and introduces local variations in hopping connectivity.
@@
We show a geometrically controlled SDE by changing the kink angle $\theta$ and map the diode efficiency by scanning the angle and the external magnetic field perpendicular to the RSOC plane. We demonstarte that the breaking of time-reversal and inversion symmetry are crucial ingredients for the manifestation of the SDE. Interestingly the maximal efficiency is obtained on the verge of the topological phase where non-trivial zero-energy bound states emerge. Beyond the non-reciprocal behavior of the supercurrent, we demonstrate the evolution of Andreev bound states at the Fermi level from trivial to non-trivial Majorana bound states (MBSs) with unique spectral characteristics. The organization of the paper is as follows: In Sec.II, we introduce the model and formalism. In Sec. III, we discuss the rich behavior of the supercurrent as a function of the kink angle and applied magnetic field, focusing on the supercurrent diode effect (SDE). Finally, in Sec. IV, we present our conclusions and provide an outlook.
@@ II Model and formalism
We consider a kinked nanostrip superconducting junction on a two-dimensional gas with Rashba interaction and variable kink angle. The superconducting leads are considered at a relative angle $\theta$ with respect to the positive orientation of the $x$-axis of the square lattice (see Fig. 1). The region defined by the junction determines the system domain $\mathcal{D}$, with its area given by $A=L_{W}(2L_{S}+L_{N})$. The lengths of the superconducting and normal regions projected along the $x$-axis are denoted by $L_{S}$ and $L_{N}$ respectively, while $L_{W}$ represents the width of the junction (see supplemental materials at VI). For every kinked realization $\theta$ the system experiences random surface roughness where some neighboring sites are connected by hopping and others are not (see Fig. 1(b)). The latter effects simulate the detrimental impact of lattice disorder, which is extremely relevant in experiments.
Figure 2: (a) Phase diagram of the critical current as a function of $\theta$ and $M_{z}$, the white lines correspond to $\theta_{n}=\arctan(3/n)$ with $n=4,6,8,10,12,18,24$. (b) Selected horizontal cuts from panel (a) showing a diode effect strongly marked for high values of $\theta$. (c) Phase diagram of the diode efficiency $\eta$ and (d) field-optimization of $\eta$ as a function of $\theta$. The superimposed orange line depicts the moving average computed over a 25-point data window. (e) Phase diagram of the zero-phase current behavior, with select horizontal cuts provided in panel (f). Notably, red regions correspond to the nucleation of $\pi/2$-junctions, easily recognizable by the enhanced regions in the magenta and dark magenta curves of the panel (f) ($\theta=0.4$, $0.7$). The $\theta$ values chosen in panels (b) and (f) are also signaled in the phase diagrams (a) and (e) by means of horizontal dashed lines.
@@ III Results
Figure 3: Classification of the sub-gap states, ABSs, MBSs and ZABSs, by means of the CPRs (a), (d), (g), the probability densities $|\psi_{i}^{2}|$ (b), (e), (h) and spectra and DOSs (c), (f), (i). The graphs and the DOSs are calculated respectively by fixing $\phi$ at the value marked by the red point in the CPRs and the vertical red cut in the spectra. The graphs reproduce the geometry of a nano-strip with $L_{S}=100$, $L_{N}=3$, $L_{W}=3$, being the radius of the circles proportional to $|\psi_{i}^{2}|$ ($i=1,\dots,L$). The spectral energy $E$ and the Majorana polarization $P_{M}(E)$ of these states are also reported. For the current and energy plots we fix respectively $et/\hbar=1$ and $t=1$.
@@
Apart from MBSs, the peculiar intertwining between the effects controlled by the geometry (kink angle) and the magnetic energy engenders additional sub-gap quasiparticle modes. We classify all the sub-gap states into three categories, distinguished by their spectral properties and influence on the CPR (Fig. 3). Trivial Andreev bound states (ABSs) are associated with sinusoidal CPRs (Fig. 3(a)) and their probability densities are spread over the whole system size (Fig. 3(b)), carrying a sub-gap excitation energy well above the Fermi level (Fig. 3(g)). These states can be found in the region where $I_{c}$ decreases with respect to $M_{z}$. MBSs, on the other hand, are associated with an enhancement of $I_{c}$ vs $M_{z}$ curves and exhibit a sawtooth CPR with a single nodal point in the phase-periodicity range (Fig. 3(d)). They are localized at the four edges of the junction (Fig. 3(e)) and reside at the Fermi energy (Fig. 3(f)) [71, 72]. Finally, Zero-energy Andreev bound states (ZABSs) can be found in the region where $I_{c}\sim I_{c}(M_{z}=0)$, before the critical current increases. The phase-space region, defined by $(\theta,M_{z})$ and hosting ZABSs, exhibits complex boundaries, the prediction of which is a rather complex task. Concerning their influence on the transport properties, ZABSs exhibit a double-sawtooth CPR (Fig. 3(g)) and, compared to MBSs, exhibit localization on a smaller number of lattice sites, placed at the left corner of the superconducting leads (Fig. 3(h)). They also share the same energy of MBSs for two specific values of the superconducting phase difference (Fig. 3(i)). However, despite their similarities to MBSs, neither ABSs nor ZABSs carry topological charge, as evidenced by the Majorana polarization ($P_{M}$[73, 74], which is indicated in the insets of panels (b), (e), and (f). Indeed, $P_{M}$ is a measure of topological quasiparticle weight in the Nambu space. The local Majorana polarization is a function of lattice sites ($i$, $j$ $\in\mathcal{D}$) and energy: $P^{loc}_{M}(i,j,\omega)=\sum_{n}\bigl{(}\sum_{\sigma}u_{n,i,j,\sigma}v_{n,i,j,\sigma}\bigr{)}(\delta(\omega-E_{n})+\delta(\omega+E_{n}))$, with $u$ and $v$ the particle and hole components of the Bogoliubov wavefunction, while $\sigma$ is the spin. Let us note that, by choosing $\omega=0$, the total Majorana polarization $P_{M}=|\sum_{i=1}^{L_{S}/4}\sum_{j=1}^{L_{w}}P^{loc}_{M}(i,j,0)|$ is equal to $1$ for genuine Majorana fermions while $P_{M}\ll 1$ for ABSs and ZABSs. Moreover, ZABs can be understood as defect states dressed by the superconducting pairing, which originates from connectivity defects or quasi-periodicity of the lattice structure [75, 76]. $P^{loc}_{M}(i,j,0)$ defines a vector with $P^{loc}_{M,x}(i,j)=Re[P^{loc}_{M}(i,j,0)]$ and $P^{loc}_{M,y}(i,j)=Im[P^{loc}_{M},(i,j,0)]$, and both $P^{loc}_{M,x}$ and $P^{loc}_{M,y}$ are peaked functions at the system edges, i.e., they show MBSs with opposite topological charge.
@@ IV Conclusions and outlooks
In summary, we have reported the emergence of the superconducting diode effect in geometry-controlled Josephson junctions based on Rashba nanostrips. By tuning parameters such as the magnetic field and the kink angle, $M_{Z}$ and $\theta$, respectively, we achieved control over critical current and diode efficiency, which is pivotal for device optimization. Notably, we achieved remarkable optimization of diode efficiency, with values exceeding $40\%$ and a minimum efficiency of $20\%$ for $\theta\geq 0.1$. The intricate interplay between effects controlled by $\theta$ and $M_{Z}$ reveals additional intriguing phenomena, including the crossover from conventional to anomalous Josephson effects and the effective tuning of the quasiparticle spectrum. Our investigation identifies Andreev bound states, zero-energy Andreev bound states, and Majorana bound states as significant sub-gap excitations, each associated with distinct transport regimes of the junction. These phases are highly sensitive to the nanochannel bending, thus providing a relevant resource to optimize the SDE. Among the experimental platforms, the oxide interfaces between LAO and STO are promising candidates to realize the kinked nano-strip. Here it’s possible to realize the kinked junctions with typical Rashba interaction of the order of magnitude considered in our manuscript. Moreover, different techniques are experimentally available to engineer the junctions with desired geometry, like the energy ion irradiation technique [59, 77, 78] or e-beam lithography with lift off technique [37]. Another suitable platform is represented by the flakes of $2H$-$NbSe_{2}$, a layered two-dimensional (2D) transition metal dichalcogenide, a prototypical type-II superconductor that exhibits strong anisotropic responses to external magnetic field [7]. For real systems, for which the fine geometrical tuning may not be achieved with a large enough resolution, we demonstrated the robustness of the SDE by increasing the kink angle, i.e., the higher $\theta$, the larger the SDE. Moreover, we cannot exclude that actual physical systems do not present asymmetries, angular misalignments, and/or irregularities at the edges. Indeed, the presence of impurities has been demonstrated to induce transition to $\pi$-junctions in spin-orbit magnetic devices [79, 80]. This aspect suggests the potential coexistence of different phases in a realistic experimental setup close to specific junction configurations, which can be further investigated. In conclusion, our findings represent a significant advancement in both the fundamental understanding and practical engineering of Rashba-based Josephson junctions for future nanoelectronic devices.
@@ VI.3 Further results on the anomalous Josephson effect
In Fig. 5(a) we consider a horizontal cut extracted from the phase diagram depicted in Fig. 2(b), with $\theta=0.2$. A dip-peak structure is evident around $M_{z}\sim 0.055$ (vertical dashed red line), marking the critical threshold for the onset of topological phases. We observe that the spectral behavior, initially lifted from zero energy at lower fields with a sinusoidal CPR (see Fig. 5(b)), converges towards zero energy for $M_{z}\geq 0.055$ (see Figs. 5(c)-(d)), accompanied by the emergence of a sawtooth CPR in Fig. 5(d). The grey and black curves in the spectra represent the BdG sub-gap eigenenergies. The lowest energy states (black ones) are insensitive to the phase difference $\phi$ in the topological phase (Fig. 5(d)), the latter being a signature of Majorana character.